ShortyShelbyj:

What is the sum of the first seven terms of the geometric series 3 − 6 + 12 − . . . ? Hint: cap s sub n equals start fraction a sub one left parenthesis one minus r to the power of n end power right parenthesis over one minus r end fraction comma r ≠ 1, where a1 is the first term and r is the common ratio. S7 = −63 S7 = −9 S7 = 33 S7 = 129

OliveSpice89:

Ok so what do you believe it is?

snowflake0531:

This is adding/subtracting the double of the previous number 3-6+12-24+48-96+192

snowflake0531:

So add them together to get your answer

surjithayer:

\[first~term=a _{1}=3\] first we need to find the common ratio. \[common~ratio~r=\frac{ a _{n} }{ a_{n-1} }=\frac{ any~term }{ previous ~term}=\frac{ a_{2} }{ a_{1}}=\frac{ -6 }{ 3}=-2 \] here we took 2nd and first term. \[S_{n}=a_{1}\frac{ 1-r^n }{ 1-r } (for~G.P.)\] here n=7,plug the values and find the sum of first 7 terms.